# Quantum Algorithms for the Pathwise Lasso

### v1

@@ 5 Bounds in noisy regime
The results from this section are standard in the Lasso literature [Wai19, BvdG11] for the case when the regularisation path from LARS algorithm exactly minimises the Lasso cost function. We generalise them for case when $\mathcal{P}$ is an approximate regularisation path with error $\lambda\epsilon\|\widetilde{\bm{\beta}}\|_{1}$.
@@ 5.1 Slow rates
from which we get the following inequalities
$\frac{1}{2}{\|}\mathbf{X}(\bm{\beta}^{*}-\widetilde{\bm{\beta}}){\|}_{2}^{2}+\lambda(1-\epsilon){\|}\widetilde{\bm{\beta}}{\|}_{1}+\frac{1}{2}{\|}\mathbf{w}{\|}_{2}^{2}+\mathbf{w}^{\top}\mathbf{X}(\bm{\beta}^{*}-\widetilde{\bm{\beta}})$ $\leq\frac{1}{2}{\|}\mathbf{w}{\|}_{2}^{2}+\lambda{\|}\bm{\beta}^{*}{\|}_{1}\implies$
$\frac{1}{2}{\|}\mathbf{X}(\bm{\beta}^{*}-\widetilde{\bm{\beta}}){\|}_{2}^{2}+\lambda(1-\epsilon){\|}\widetilde{\bm{\beta}}{\|}_{1}+\mathbf{w}^{\top}\mathbf{X}(\bm{\beta}^{*}-\widetilde{\bm{\beta}})$ $\leq\lambda{\|}\bm{\beta}^{*}{\|}_{1}$
@@
$\frac{1}{2}{\|}\mathbf{X}(\bm{\beta}^{*}-\widetilde{\bm{\beta}}){\|}_{2}^{2}$ $\leq\mathbf{w}^{\top}\mathbf{X}(\widetilde{\bm{\beta}}-\bm{\beta}^{*})+\lambda{\|}\bm{\beta}^{*}{\|}_{1}-\lambda(1-\epsilon)\|\widetilde{\bm{\beta}}\|_{1}$
$\leq\|\mathbf{X}^{\top}\mathbf{w}\|_{\infty}\|\widetilde{\bm{\beta}}-\bm{\beta}^{*}\|_{1}+\lambda{\|}\bm{\beta}^{*}{\|}_{1}-\lambda(1-\epsilon)\|\widetilde{\bm{\beta}}\|_{1}$
$\leq\|\mathbf{X}^{\top}\mathbf{w}\|_{\infty}(\|\widetilde{\bm{\beta}}\|+\|\bm{\beta}^{*}\|_{1})+\lambda{\|}\bm{\beta}^{*}{\|}_{1}-\lambda(1-\epsilon)\|\widetilde{\bm{\beta}}\|_{1}$
@@
$\leq{\|}{\bm{\beta}}^{*}{\|}(\|\mathbf{X}^{\top}\mathbf{w}\|_{\infty}+\lambda)$
$\leq(2-\epsilon)\lambda{\|}\bm{\beta}^{*}{\|},$
where the second inequality follows from Hölder’s inequality and the third follows from the triangle inequality. ∎
@@
In the classical linear Gaussian model, the design matrix is deterministic, meaning that $\mathbf{X}\in\mathbb{R}^{n\times d}$ is fixed, while the noise vector $\mathbf{w}\in\mathbb{R}^{n}$ has i.i.d. $\mathcal{N}(0,\sigma^{2})$ entries. We then bound the probability that $\lambda\geq\|\mathbf{X}^{\top}\mathbf{w}\|_{\infty}/(1-\epsilon)$. Using that $\mathbf{X}_{i}^{\top}\mathbf{w}$ is a sub-Gaussian random variable with parameter $\sigma\|\mathbf{X}_{i}\|_{2}\leq\sigma\sqrt{Cn}$ (Fact 6), we bound the complement of this probability:
@@
In compressed sensing, it is common for the design matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ to be chosen by the user, and a standard choice is the standard Gaussian matrix with i.i.d. $\mathcal{N}(0,1)$ entries. Assume further that the noise vector $\mathbf{w}\in\mathbb{R}^{n}$ is deterministic with $\|\mathbf{w}\|_{\infty}\leq\sigma$. Thus $\mathbf{X}_{i}^{\top}\mathbf{w}$ is a sub-Gaussian random variable with parameter $\|\mathbf{w}\|_{2}\leq\sqrt{n}\|\mathbf{w}\|_{\infty}\leq\sqrt{n}\sigma$ (Fact 6). By following the exact same steps as the previous example (this time with $C=1$), we conclude that, with probability at least $1-\delta$,
@@ 5.2 Fast Rates
The proof has many variations, and we give the most direct one below. Note that the above implies
@@
$0\leq 2{\|}\mathbf{X}\bm{\Delta}{\|}_{2}^{2}$ $\leq 4\mathbf{w}^{\top}\mathbf{X}(\widetilde{\bm{\beta}}-\bm{\beta}^{*})+4\lambda{\|}\bm{\beta}^{*}{\|}_{1}-4\lambda(1-\epsilon)\|\widetilde{\bm{\beta}}\|_{1}$
$\leq 4\|\mathbf{X}^{\top}\mathbf{w}\|_{\infty}\|\bm{\Delta}\|_{1}+4\lambda{\|}\bm{\beta}^{*}{\|}_{1}-4\lambda(1-\epsilon)\|\widetilde{\bm{\beta}}\|_{1}$
$\leq\lambda\|\bm{\Delta}\|_{1}+4\lambda{\|}\bm{\beta}^{*}{\|}_{1}-4\lambda(1-\epsilon)\|\widetilde{\bm{\beta}}\|_{1}$
$\leq\lambda(\|\widetilde{\bm{\beta}}\|_{1}+\|\bm{\beta}^{\ast}\|_{1})+4\lambda{\|}\bm{\beta}^{*}{\|}_{1}-4\lambda(1-\epsilon)\|\widetilde{\bm{\beta}}\|_{1},$
from which we infer that $\|\widetilde{\bm{\beta}}\|_{1}\leq\frac{5}{3-4\epsilon}\|\bm{\beta}^{\ast}\|_{1}\leq\frac{5}{2}\|\bm{\beta}^{\ast}\|_{1}$. Let us now return to the inequality $2{\|}\mathbf{X}\bm{\Delta}{\|}_{2}^{2}\leq\lambda\|\bm{\Delta}\|_{1}+4\lambda({\|}\bm{\beta}^{*}{\|}_{1}-\|\widetilde{\bm{\beta}}\|_{1})+4\lambda\epsilon\|\widetilde{\bm{\beta}}\|_{1}$ and further decompose the right-hand side into the respective partitioning sets $S$ and $S^{c}$ by using that $\bm{\beta}_{S}^{*}+\bm{\Delta}_{S}=\widetilde{\bm{\beta}}_{S}$ and $\bm{\Delta}_{S^{c}}=\widetilde{\bm{\beta}}_{S^{c}}$, since $\bm{\beta}_{S^{c}}^{*}=0$. We also use the inequality $\|\widetilde{\bm{\beta}}\|_{1}\leq\frac{5}{2}\|\bm{\beta}^{\ast}\|_{1}=\frac{5}{2}\|\bm{\beta}^{\ast}_{S}\|_{1}$. Therefore
@@
$\leq\lambda(\|\bm{\Delta}_{S}\|_{1}+\|\bm{\Delta}_{S^{c}}\|_{1})+4\lambda({\|}\bm{\beta}^{*}{\|}_{1}-\|\widetilde{\bm{\beta}}\|_{1})+10\lambda\epsilon\|\bm{\beta}^{\ast}_{S}\|_{1}$
@@
$\overset{(\ast)}{\leq}\lambda(\|\bm{\Delta}_{S}\|_{1}+\|\bm{\Delta}_{S^{c}}\|_{1})+4\lambda(\|\bm{\Delta}_{S}\|_{1}-\|\bm{\Delta}_{S^{c}}\|_{1})+10\lambda\epsilon\|\bm{\beta}^{\ast}_{S}\|_{1}$
@@
where the inequality ()($\ast$)( ∗ ) follows from
$\|\bm{\beta}_{S}^{*}\|_{1}-\|\bm{\beta}_{S}^{*}+\bm{\Delta}_{S}\|_{1}\leq\|\bm{\beta}_{S}^{*}\|_{1}-(\|\bm{\beta}_{S}^{*}\|_{1}-\|\bm{\Delta}_{S}\|_{1})=\|\bm{\Delta}_{S}\|_{1}.$
@@
In the first case, we find that $\bm{\Delta}\in\mathbb{C}_{5}(S)$, which means that we can now apply the RE condition. More specifically, from $2{\|}\mathbf{X}\bm{\Delta}{\|}_{2}^{2}+3\lambda\|\bm{\Delta}_{S^{c}}\|_{1}\leq 15\lambda{\|}\bm{\Delta}_{S}{\|}_{1}$ we get
$2{\|}\mathbf{X}\bm{\Delta}{\|}_{2}^{2}+3\lambda\|\bm{\Delta}\|_{1}$ $\leq 18\lambda{\|}\bm{\Delta}_{S}{\|}_{1}\leq 18\lambda\sqrt{|S|}\|\bm{\Delta}\|_{2}\leq 18\lambda\sqrt{\frac{|S|}{n\kappa}}\|\mathbf{X}\bm{\Delta}\|_{2}\leq 81\lambda^{2}\frac{|S|}{n\kappa}+\|\mathbf{X}\bm{\Delta}\|_{2}^{2},$
@@ 6 Discussion and future work
We studied and quantised the LARS pathwise algorithm (or homotopy method) proposed by Efron et al. [EHJT04] and Osborne et al. [OPT00b, OPT00a] which produces the set of Lasso solutions for varying the penalty term $\lambda$. By assuming quantum access to the Lasso input, we proposed two quantum algorithms. The first one (Algorithm 1) simply replaces the classical search within the joining time calculation with the quantum minimum-finding subroutine from Dürr and Høyer [DH96]. Similar to the classical LARS algorithm, it outputs the exact Lasso path, but has an improved runtime by a quadratic factor in the number of features $d$. Our second quantum algorithm (Algorithm 3), on the other hand, outputs an approximate Lasso path $\widetilde{\mathcal{P}}$ by computing the joining times up to some error. This is done by approximating the joining times $\Lambda_{i}^{\rm join}(\mathcal{A})=\frac{\mathbf{X}_{i}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu})}{\pm 1-\mathbf{X}_{i}^{\top}\mathbf{X}_{\mathcal{A}}\bm{\theta}}$ using quantum amplitude estimation [BHMT02] and finding their maximum via the approximate quantum minimum-finding subroutine from Chen and de Wolf [CdW23]. Consequently, the runtime of our approximate quantum LARS algorithm is quadratically improved in both the number of features $d$ and observations $n$. We stabilised the correctness of Algorithm 3 by employing an approximate version of the KKT conditions and a duality gap, and showed that $\widetilde{\mathcal{P}}$ is a minimiser of the Lasso cost function up to additive error $\lambda\epsilon\|\widetilde{\bm{\beta}}(\lambda)\|_{1}$. Finally, we dequantised Algorithm 3 and proposed an approximate classical LARS algorithm (Algorithm 4) based on sampling that retains the quadratic improvement in the number of observations $n$.
The time complexity of our approximate LARS algorithms, both classical and quantum, directly depends on the design matrix $\mathbf{X}$ via the quantity $\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}\|_{2}$, the mutual incoherence, and the mutual overlap between the design matrix and the vector of observations $\mathbf{y}$. If $\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}\|_{2}$ is a constant and mutual incoherence and overlap are both bounded away from 00, then the approximate quantum LARS algorithm’s complexity is $\widetilde{O}(\sqrt{n|\mathcal{I}|}/\epsilon+n|\mathcal{A}|+|\mathcal{A}|^{2})$ per iteration. For small sizes of the active set, $|\mathcal{A}|=O(1)$, we obtain the overall quadratic improvement $\widetilde{O}(\sqrt{nd})$ over the classical LARS algorithm. However, for large sizes as $|\mathcal{A}|=\Omega(n)$, it is more advantageous to employ the simple quantum LARS algorithm since its complexity is just $\widetilde{O}(n\sqrt{d})$. The simple quantum LARS algorithm, moreover, does not depend on $\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}\|_{2}$ and the mutual incoherence and overlap, and therefore should be superior for a wider class of design matrices. We proved, nonetheless, that Algorithm 3 exhibits the complexity $\widetilde{O}(\sqrt{nd})$ per iteration for the case when $\mathbf{X}$ is a standard Gaussian random matrix with high probability.
@@
We point out some future directions of research following our work. In Algorithms 3 and 4, the crossing times are computed exactly and the joining times are estimated up to some error. A natural extension would be to allow errors also in computing the crossing times. This would require new techniques in order to retain the path solution continuity. Another direction is to reduce the number of iterations of the LARS algorithm. In the classical setting, Mairal and Yu [MY12] proposed an approximate LARS algorithm with a maximum number of iterations by employing first-order optimisation methods when two kinks are too close to each other. Finally, it would be interesting to design a fully quantum LARS algorithm by using efficient quantum subroutines for matrix multiplication and matrix inversion, e.g. based on block-encoding techniques [LC19, GSLW19, CGJ19].
@@
Our algorithms are designed to work only in the fault tolerant regime. Another future direction is to study the LARS algorithm in the NISQ (Noisy Intermediate-Scale Quantum) setting [Pre18], where algorithms operate on a relatively small number of qubits and have shallow circuit depths. In addition, while we have studied the plain-vanilla LARS algorithm, other Lasso related algorithms still remain unexplored in the quantum setting. For example, fused Lasso [${}^{+}$05, ${}^{+}$16, Hoe10], which penalises the $\ell_{1}$-norm of both the coefficients and their successive differences, is a generalisation of Lasso with applications to support vector classifier [Gun98, RML14, SV99]. On the other hand, grouped Lasso is a generalised model for linear regression with $\ell_{1}$ and $\ell_{2}$-penalties [SFHT13, FHT10b]. This model is used in settings where the design matrix can be decomposed into groups of submatrices. Simon et al. [SFHT13] proposed a sparse grouped Lasso algorithm which finds applications in sparse graphical modelling. We believe that similar techniques employed in this work could be applied to these alternative Lasso settings.

### v2

@@ 5 Bounds in noisy regime
The results from this section are standard in the Lasso literature [114, 9] for the case when the regularisation path from LARS algorithm exactly minimises the Lasso cost function. We generalise them for case when $\widetilde{\mathcal{P}}$ is an approximate regularisation path with error $\lambda\epsilon\|\widetilde{\bm{\beta}}\|_{1}$.
@@ 5.1 Slow rates
from which we get the following inequality
$\frac{1}{2}{\|}\mathbf{X}(\bm{\beta}^{*}-\widetilde{\bm{\beta}}){\|}_{2}^{2}+\lambda(1-\epsilon){\|}\widetilde{\bm{\beta}}{\|}_{1}+\frac{1}{2}{\|}\mathbf{w}{\|}_{2}^{2}+\mathbf{w}^{\top}\mathbf{X}(\bm{\beta}^{*}-\widetilde{\bm{\beta}})$ $\leq\frac{1}{2}{\|}\mathbf{w}{\|}_{2}^{2}+\lambda{\|}\bm{\beta}^{*}{\|}_{1}\iff$
$\frac{1}{2}{\|}\mathbf{X}(\bm{\beta}^{*}-\widetilde{\bm{\beta}}){\|}_{2}^{2}+\lambda(1-\epsilon){\|}\widetilde{\bm{\beta}}{\|}_{1}+\mathbf{w}^{\top}\mathbf{X}(\bm{\beta}^{*}-\widetilde{\bm{\beta}})$ $\leq\lambda{\|}\bm{\beta}^{*}{\|}_{1}$ (17)
@@
$\frac{1}{2}{\|}\mathbf{X}(\bm{\beta}^{*}-\widetilde{\bm{\beta}}){\|}_{2}^{2}$ $\leq\mathbf{w}^{\top}\mathbf{X}(\widetilde{\bm{\beta}}-\bm{\beta}^{*})+\lambda{\|}\bm{\beta}^{*}{\|}_{1}-\lambda(1-\epsilon)\|\widetilde{\bm{\beta}}\|_{1}$ (Equation 17)
$\leq\|\mathbf{X}^{\top}\mathbf{w}\|_{\infty}\|\widetilde{\bm{\beta}}-\bm{\beta}^{*}\|_{1}+\lambda{\|}\bm{\beta}^{*}{\|}_{1}-\lambda(1-\epsilon)\|\widetilde{\bm{\beta}}\|_{1}$ (Hölder’s inequality)
$\leq\|\mathbf{X}^{\top}\mathbf{w}\|_{\infty}(\|\widetilde{\bm{\beta}}\|+\|\bm{\beta}^{*}\|_{1})+\lambda{\|}\bm{\beta}^{*}{\|}_{1}-\lambda(1-\epsilon)\|\widetilde{\bm{\beta}}\|_{1}$ (Triangle inequality)
@@
$\leq{\|}{\bm{\beta}}^{*}{\|}(\|\mathbf{X}^{\top}\mathbf{w}\|_{\infty}+\lambda)$ ($\|\mathbf{X}^{\top}\mathbf{w}\|_{\infty}\leq(1-\epsilon)\lambda$)
$\leq(2-\epsilon)\lambda{\|}\bm{\beta}^{*}{\|}.$ ($\|\mathbf{X}^{\top}\mathbf{w}\|_{\infty}\leq(1-\epsilon)\lambda$)  ∎
@@
In the classical linear Gaussian model, the design matrix is deterministic, meaning that $\mathbf{X}\in\mathbb{R}^{n\times d}$ is fixed, while the noise vector $\mathbf{w}\in\mathbb{R}^{n}$ has i.i.d. $\mathcal{N}(0,\sigma^{2})$ entries. We then bound the probability that $\lambda\geq\|\mathbf{X}^{\top}\mathbf{w}\|_{\infty}/(1-\epsilon)$. Using that $\mathbf{X}_{i}^{\top}\mathbf{w}$ is a sub-Gaussian random variable with parameter $\sigma\|\mathbf{X}_{i}\|_{2}\leq\sigma\sqrt{Cn}$ (Fact 7), we bound the complement of this probability:
@@
In compressed sensing, it is common for the design matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ to be chosen by the user, and a standard choice is the standard Gaussian matrix with i.i.d. $\mathcal{N}(0,1)$ entries. Assume further that the noise vector $\mathbf{w}\in\mathbb{R}^{n}$ is deterministic with $\|\mathbf{w}\|_{\infty}\leq\sigma$. Thus $\mathbf{X}_{i}^{\top}\mathbf{w}$ is a sub-Gaussian random variable with parameter $\|\mathbf{w}\|_{2}\leq\sqrt{n}\|\mathbf{w}\|_{\infty}\leq\sqrt{n}\sigma$ (Fact 7). By following the exact same steps as the previous example (this time with $C=1$), we conclude that, with probability at least $1-\delta$,
@@ 5.2 Fast Rates
The proof has many variations, and we give the most direct one below. Note the above implies
@@
$0\leq 2{\|}\mathbf{X}\bm{\Delta}{\|}_{2}^{2}$ $\leq 4\mathbf{w}^{\top}\mathbf{X}(\widetilde{\bm{\beta}}-\bm{\beta}^{*})+4\lambda{\|}\bm{\beta}^{*}{\|}_{1}-4\lambda(1-\epsilon)\|\widetilde{\bm{\beta}}\|_{1}$ (Equation 17)
$\leq 4\|\mathbf{X}^{\top}\mathbf{w}\|_{\infty}\|\bm{\Delta}\|_{1}+4\lambda{\|}\bm{\beta}^{*}{\|}_{1}-4\lambda(1-\epsilon)\|\widetilde{\bm{\beta}}\|_{1}$ (Hölder’s inequality)
$\leq\lambda\|\bm{\Delta}\|_{1}+4\lambda{\|}\bm{\beta}^{*}{\|}_{1}-4\lambda(1-\epsilon)\|\widetilde{\bm{\beta}}\|_{1}$ ($4\|\mathbf{X}^{\top}\bm{\Delta}\|_{\infty}\leq\lambda$)
$\leq\lambda(\|\widetilde{\bm{\beta}}\|_{1}+\|\bm{\beta}^{\ast}\|_{1})+4\lambda{\|}\bm{\beta}^{*}{\|}_{1}-4\lambda(1-\epsilon)\|\widetilde{\bm{\beta}}\|_{1},$ (Triangle inequality)
from which we infer that $\|\widetilde{\bm{\beta}}\|_{1}\leq\frac{5}{3-4\epsilon}\|\bm{\beta}^{\ast}\|_{1}\leq\frac{5}{2}\|\bm{\beta}^{\ast}\|_{1}$ since $\epsilon\leq\frac{1}{4}$. Let us now return to the inequality $2{\|}\mathbf{X}\bm{\Delta}{\|}_{2}^{2}\leq\lambda\|\bm{\Delta}\|_{1}+4\lambda({\|}\bm{\beta}^{*}{\|}_{1}-\|\widetilde{\bm{\beta}}\|_{1})+4\lambda\epsilon\|\widetilde{\bm{\beta}}\|_{1}$ and further decompose the right-hand side into the respective partitioning sets $S$ and $S^{c}$ by using that $\bm{\beta}_{S}^{*}+\bm{\Delta}_{S}=\widetilde{\bm{\beta}}_{S}$ and $\bm{\Delta}_{S^{c}}=\widetilde{\bm{\beta}}_{S^{c}}$, since $\bm{\beta}_{S^{c}}^{*}=\mathbf{0}$. We also use the inequality $\|\widetilde{\bm{\beta}}\|_{1}\leq\frac{5}{2}\|\bm{\beta}^{\ast}\|_{1}=\frac{5}{2}\|\bm{\beta}^{\ast}_{S}\|_{1}$. Therefore
@@
$=\lambda(\|\bm{\Delta}_{S}\|_{1}+\|\bm{\Delta}_{S^{c}}\|_{1})+4\lambda({\|}\bm{\beta}^{*}{\|}_{1}-\|\widetilde{\bm{\beta}}\|_{1})+10\lambda\epsilon\|\bm{\beta}^{\ast}_{S}\|_{1}$
@@
$\leq\lambda(\|\bm{\Delta}_{S}\|_{1}+\|\bm{\Delta}_{S^{c}}\|_{1})+4\lambda(\|\bm{\Delta}_{S}\|_{1}-\|\bm{\Delta}_{S^{c}}\|_{1})+10\lambda\epsilon\|\bm{\beta}^{\ast}_{S}\|_{1}$ (Equation 18)
@@
where we used that
$\|\bm{\beta}_{S}^{*}\|_{1}-\|\bm{\beta}_{S}^{*}+\bm{\Delta}_{S}\|_{1}\leq\|\bm{\beta}_{S}^{*}\|_{1}-(\|\bm{\beta}_{S}^{*}\|_{1}-\|\bm{\Delta}_{S}\|_{1})=\|\bm{\Delta}_{S}\|_{1}.$ (18)
@@
In the first case, we find that $\|\bm{\Delta}_{S^{c}}\|_{1}\leq 5\|\bm{\Delta}_{S}\|_{1}\implies\bm{\Delta}\in\mathbb{C}_{5}(S)$, which means that we can now apply the RE condition. More specifically, from $2{\|}\mathbf{X}\bm{\Delta}{\|}_{2}^{2}+3\lambda\|\bm{\Delta}_{S^{c}}\|_{1}\leq 15\lambda{\|}\bm{\Delta}_{S}{\|}_{1}$ we get
$2{\|}\mathbf{X}\bm{\Delta}{\|}_{2}^{2}+3\lambda\|\bm{\Delta}\|_{1}$ $\leq 18\lambda{\|}\bm{\Delta}_{S}{\|}_{1}\leq 18\lambda\sqrt{|S|}\|\bm{\Delta}\|_{2}\overset{\rm RE}{\leq}18\lambda\sqrt{\frac{|S|}{n\kappa}}\|\mathbf{X}\bm{\Delta}\|_{2}\leq 81\lambda^{2}\frac{|S|}{n\kappa}+\|\mathbf{X}\bm{\Delta}\|_{2}^{2},$
@@ 6 Discussion and future work
We studied and quantised the LARS pathwise algorithm (or homotopy method) proposed by Efron et al. [32] and Osborne et al. [79, 80] which produces the set of Lasso solutions for varying the penalty term $\lambda$. By assuming quantum access to the Lasso input, we proposed two quantum algorithms. The first one (Algorithm 2) simply replaces the classical search within the joining time calculation with the quantum minimum-finding subroutine from Dürr and Høyer [25]. Similar to the classical LARS algorithm, it outputs the exact Lasso path, but has an improved runtime by a quadratic factor in the number of features $d$. Our second quantum algorithm (Algorithm 3), on the other hand, outputs an approximate Lasso path $\widetilde{\mathcal{P}}$ by computing the joining times up to some error. This is done by approximating the joining times using quantum amplitude estimation [10] and finding their maximum via the approximate quantum minimum-finding subroutine from Chen and de Wolf [23]. Consequently, the runtime of our approximate quantum LARS algorithm is quadratically improved in both the number of features $d$ and observations $n$. We established the correctness of Algorithm 3 by employing an approximate version of the KKT conditions and a duality gap, and showed that $\widetilde{\mathcal{P}}$ is a minimiser of the Lasso cost function up to additive error $\lambda\epsilon\|\widetilde{\bm{\beta}}(\lambda)\|_{1}$. Finally, we dequantised Algorithm 3 and proposed an approximate classical LARS algorithm (Algorithm 4) based on sampling.
The time complexity of our approximate LARS algorithms, both classical and quantum, directly depends on the design matrix $\mathbf{X}$ via the quantity $\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}_{\mathcal{A}}\|_{2}$, the mutual incoherence of $\mathbf{X}$, and the mutual overlap between $\mathbf{X}$ and the vector of observations $\mathbf{y}$. When $\mathbf{X}$ is a random matrix from the $\mathbf{\Sigma}$-Gaussian ensemble, where $\mathbf{\Sigma}$ is a covariance matrix with “good” properties, and $\mathbf{y}$ is sparse, we showed that these three quantities are well bounded and that Algorithm 3 and Algorithm 4 have complexities $\widetilde{O}(\sqrt{d}/\epsilon)$ and $\widetilde{O}(d/\epsilon^{2})$ per iteration, respectively, which exponentially improves the dependence on $n$ compared to the $O(nd)$ complexity of the standard LARS algorithm. The condition that $\mathbf{y}$ is sparse follows naturally from the fact that there could be data situations in which $\mathbf{X}$ being Gaussian leads to very small, in absolute magnitude, responses in the dependent variable $\mathbf{y}$. Another reason for $\mathbf{y}$ being sparse is inaccurate or incomplete measurements [20]. Usually, sparsity can be artificially generated by the corresponding matrix $\mathbf{X}$ through setting some of the regression coefficients in $\bm{\beta}^{\ast}$ to zero or setting some of the components in $\mathbf{y}$ to zero. In both cases, our algorithms would be robust enough and exhibit speed up to recover the approximate original coefficient vector. Even though we only proved such result for random Gaussian matrices, we conjecture that it should also apply to other matrices.
@@
We point out some future directions of research following our work. In Algorithm 3 and Algorithm 4, the crossing times are computed exactly and the joining times are estimated up to some error. A natural extension would be to allow errors also in computing the crossing times. This would require new techniques in order to retain the path solution continuity. Another direction is to reduce the number of iterations of the LARS algorithm. In the classical setting, Mairal and Yu [78] proposed an approximate LARS algorithm with a maximum number of iterations by employing first-order optimisation methods when two kinks are too close to each other. Finally, it would be interesting to design a fully quantum LARS algorithm by using efficient quantum subroutines for matrix multiplication and matrix inversion, e.g., based on block-encoding techniques [61, 44, 16].
@@
Our algorithms are designed to work only in the fault tolerant regime. Another future direction is to study the LARS algorithm in the NISQ (Noisy Intermediate-Scale Quantum) setting [83], where algorithms operate on a relatively small number of qubits and have shallow circuit depths. In addition, while we have studied the plain-vanilla LARS algorithm, other Lasso related algorithms still remain unexplored in the quantum setting. For example, fused Lasso [107, 122, 50], which penalises the $\ell_{1}$-norm of both the coefficients and their successive differences, is a generalisation of Lasso with applications to support vector classifier [45, 87, 101]. On the other hand, grouped Lasso is a generalised model for linear regression with $\ell_{1}$ and $\ell_{2}$-penalties [94, 34]. This model is used in settings where the design matrix can be decomposed into groups of submatrices. Simon et al. [94] proposed a sparse grouped Lasso algorithm which finds applications in sparse graphical modelling. We believe that similar techniques employed in this work could be applied to these alternative Lasso settings.